OpenStudy (anonymous):

Kayla has $50. Nina has $80. In how many months will they both have the same amount of money if Kayla saves $50 per month and Nina saves $40?

OpenStudy (anonymous):

i will give a medal

OpenStudy (lochana):

kayla has lowest money. but she saves more than nina. so definitely, one day kayla will have more money than nina

OpenStudy (lochana):

say in x months, kayla has M money so M = x*$50 + $50

OpenStudy (lochana):

in x months, nina has L money L = x*$40 + 80

OpenStudy (lochana):

now lets find the value of x(months) when L = M (nina's money and kayla's money) L = M x*$40 + 80 =x*$50 + $50

OpenStudy (lochana):

now find x

OpenStudy (lochana):

and tell me the answer:)

OpenStudy (anonymous):

they gonna have to save for a while, because they my sugar mommas

OpenStudy (anonymous):

\[50 m+50=40 m+80 \] m = 3